OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. 2x2 - 28x + 98 = 0 A. {-7, 7} B. {7} C. {-2, 49} D. no real number solution Is it A ?

jhonyy9 (jhonyy9):

B.

OpenStudy (savvy):

B.

OpenStudy (anonymous):

D=b^2-4ac we get ... D=784-76048 so the D is minus and square root of minus is a complex number... so the answer is ... NO REAL NUMBER SOLUTION !!

OpenStudy (savvy):

\[2x ^{2} - 28x + 98 \] \[x^2 - 14x +49\] \[(x-7)^2 = 0\] x=7

OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. 2x2 - 28x + 98 = 0 read Carefully !!